If the pH of a 1.00-in. rainfall over 1300 miles2 is 3.40, how many kilograms of sulfuric acid, H2SO4, are present, assuming that it is the only acid contributing to the pH? For sulfuric acid, Ka1 is very large and Ka2 is 0.012. Express your answer to two significant figures and include the appropriate units.
Answer #1

The ditotiation reaction is :

H​​​​​​​​​2​​​​​​​SO​​​​​​​4(lq)-----> 2H+ ​​​​(ag) + SO​​​​​​​4​​​​​​​​​​​​​​^2​​​​​​​- (ag)

If the pH is pH= -log[H​​​​​​​​​​​​​​+] 
Then the concentration of H + is [H+] = 10^-pH 

[H+]=10^-3.40 = 3.98x10-4M

  • Now de need the liters

1.00 in =0.0254m

1300milles2 *(25998.11m2/1miles2)=33.79×106 m​​​​​2

teh volume well be

(0.0254×33.79×10^6)m3 =858,266m3. 1m3=1000 l

V= 858,266m3(1000l/1m3)

V=856.266×106 l

MolH​​​​​​​​​​​​​​​​​​​​​2SO4=2(3.98×10-4 M )*856.266×106 =681.587×103mol

Molar mass H2SO4 = 98.079g/mol

gH2SO4=681.587×103 mol (98.078g/mol)= 66.84×106g

KgH2SO4=66.84×106g(1Kg/1000g)

KgH2SO4= 66.84×10​​​​​​​4 Kg

​​​